OpenStudy (anonymous):

Suppose you bought a 100-foot fence that you wanted to put around a square section of yard that will be used as a garden. The side length of your garden is 13 feet. The fencing was on sale for $12.60 per foot. How much fencing is left over?

OpenStudy (anonymous):

fencing left = total length of fence - total perimeter of square garden = 100 - (4*13) = 100 - 52 = 48 feet lol

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

nice copy and paste! lol

OpenStudy (anonymous):

ikr :P

OpenStudy (anonymous):

thanks @zaphodplaysitsafe

OpenStudy (anonymous):

theres three parts to the question i just posted it

OpenStudy (anonymous):

ok lol